OpenStudy (anonymous):

let z1=10+6i, z2=4+6i,if z is any complex number such that the argument z-z1/z-z2=pie/4 then the value of |z-7-9i| is?

OpenStudy (anonymous):

@Sambhavvinaykya

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

take conjugation first and separate the real and imaginary part

OpenStudy (anonymous):

can u solve whole of it.difficult to solve

OpenStudy (anonymous):

hmm ok after take conjugation and you get \[\frac{ ((x-10)+i(y-6))((x-4)-i(y-6)) }{(x-4)^{2}+(y-6)^{2}}\]

OpenStudy (anonymous):

confused.

OpenStudy (asnaseer):

@danish071996 - what part are you confused on?

OpenStudy (asnaseer):

I believe what @Sambhavvinaykya did was to first let:\[z=x+iy\]and then substitue this into:\[\frac{z-z_1}{z-z_2}\]does that make sense so far?

OpenStudy (asnaseer):

@danish071996 - if you don't answer then it becomes very difficult to help

OpenStudy (anonymous):

yes i can understand that

OpenStudy (asnaseer):

ok, so if do the substitution we get:\[\frac{x+iy-(10+6i)}{x+iy-(4+6i)}=\frac{(x-10)+i(y-6)}{(x-4)+i(y-6)}\]agreed?

OpenStudy (anonymous):

yes.after that

OpenStudy (asnaseer):

what he did next was to multiply the numerator and the denominator by the complex conjugate of the denominator

OpenStudy (asnaseer):

that would leave you with a non-complex number in the denominator - understand?

OpenStudy (anonymous):

yes.bt how to find the value?

OpenStudy (asnaseer):

so, if we follow through we get:\[\frac{(x-10)+i(y-6)}{(x-4)+i(y-6)}=\frac{ ((x-10)+i(y-6))((x-4)-i(y-6)) }{(x-4)^{2}+(y-6)^{2}}\]\[=\frac{x^2-14x+y^2-12y+76+6(y-6)i}{(x-4)^2+(y-6)^2}\]

OpenStudy (asnaseer):

your question states that the argument of this complex number is \(\pi/4\) - do you know what an argument of a complex number is?

OpenStudy (anonymous):

yes i know

OpenStudy (asnaseer):

good, so you know that if we had some complex number, say \(a+ib\), and we were told that its argument was \(\pi/4\), then this implies:\[\frac{\pi}{4}=\tan^{-1}(\frac{b}{a})\]agreed?

OpenStudy (anonymous):

yes.agree

OpenStudy (asnaseer):

good, this also therefore implies that:\[\tan(\frac{\pi}{4})=\frac{b}{a}=1\]\[\therefore b=a\]agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

good, now if we look back at the result we got:\[\frac{x^2-14x+y^2-12y+76+6(y-6)i}{(x-4)^2+(y-6)^2}\]this implies that the real and imaginary components are equal. Also, since the denominator is a real number, we can just concentrate on the numerator and deduce that:\[x^2-14x+y^2-12y+76=6(y-6)\]agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so now lets rearrange this a little:\[x^2-14x+y^2-12y+76=6y-36\]\[\therefore x^2-14x+y^2-18y+112=0\]\[(x-7)^2+(y-9)^2-18=0\]\[\therefore (x-7)^2+(y-9)^2=18\]agreed?

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

now we are almost there, the question is asking you to find:\[|z-7-9i|\]if we substitute \(z=x+iy\) into this we get:\[|z-7-9i|=|x+iy-7-9i|=|(x-7)+i(y-9)|=?\]hopefully you can complete this from here?

OpenStudy (anonymous):

yes.i got it.thanks

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

sorry internet connectivity got lost so couldnt reply @danish071996 nice explanation @asnaeer

OpenStudy (akonkel):

@danish071996